OpenStudy (anonymous):

Solve the equation: for 0° ≤ x ≤ 360° cos (45° + x) + cos (45° - x) = √ 6 sin x

OpenStudy (dumbcow):

use angle sum/difference identities

OpenStudy (anonymous):

We know that : \[\cos (a+b)+\cos(a-b)=2\cos a\cos b\] So use this formula to find that : \[2\cos 45^\circ\cos x=\sqrt6\sin x\] Can you continue from here ?

OpenStudy (anonymous):

Thanks! let me disgust =)

OpenStudy (anonymous):

May I know how to get √6 sinx = 2 cos a cos b?

OpenStudy (anonymous):

Look ! The left side of the equation is : \[\cos (45° + x) + \cos (45° - x)\] which equals : \[2\cos 45^\circ\cos x\]

OpenStudy (anonymous):

oh, got it. but I am stuck at: √2 cos x = √6 sin x cos x = √3 sin x am I right so far? the the answer is 30° & 210°

OpenStudy (anonymous):

Ok ! You get : \[\sqrt3\sin x=\cos x\] bu dividing over cos x we get : \[\sqrt3\tan x=1\] then : \[\tan x=\frac1{\sqrt3}\] That means : \[x=30^\circ\]

OpenStudy (anonymous):

wao, I wish I am as smart as you! I couldn't see that "far". Thank you so much again!!

OpenStudy (anonymous):

There is no need to be smart to solve that type of exercises ! :) It needs only to know the formulas ! That's all ! Anyway, you are welcome !